# Unsupervised detection of semantic correlations in big data

### v1

@@ Abstract
Correlations among the features defining a data point force a data set to live in a manifold $\mathcal{M}$ whose geometry encodes the correlation structure of data. The most fundamental property of $\mathcal{M}$ is its intrinsic dimensionBennett [1969] which is qualitatively defined by the number of independent variables needed to parameterize it without significant information loss. If one was able to estimate the intrinsic dimension in a generic big data scenario, one would be able to quantify the strength of correlations in real-world datasets, like a corpus of sentences or a large set of measurements from a camera. However, real-world data are represented in feature spaces that are outrageously large, and most of the available methods meet serious problems if the intrinsic dimension of the manifold is of the order of hundreds or moreCamastra and Staiano [2016]. We developed an intrinsic dimension estimator that generalizes the concept of the intrinsic dimension to the case of binary variables. We show that this allows circumventing the curse of dimensionality, and estimating reliably the intrinsic dimension even when it is of order 100,000. This result paves the way to a plethora of applications, allowing to address questions which have thus far been, to the best of our knowledge, intractable. For example, we can measure the range of semantic correlations in a corpus of sentences with hundreds of words, and we demonstrate that these correlations are destroyed when shorter, unrelated sentences are joined together. We show that our approach is effective in a variety of other fields and topics, from the study of phase transitions and of frustration-induced disorder in statistical physics, to image classification.
Detecting and quantifying correlations in generic real-world data is a complex task since features live in high dimensional spaces and have interactions of arbitrary range and nature. For example, in an image, the meaning of a scene can emerge from the simultaneous presence of two objects that are physically far away. Moreover, semantic correlations are often associated with collective effects, such as the simultaneous presence of three or more specific objects. A prominent example where non-trivial correlations emerge is the feature space of deep neural networks. Those spaces typically lack a well-defined order, and neurons that are close in the architecture are not necessarily coding related information. This hinders the direct interpretability of those feature spaces, leading many scientists, even today, to consider neural networks as "black boxes".
@@
In order to estimate the ID in very large feature spaces, we here renounce to develop a general-purpose ID estimator, and we focus on features that are binary or can be approximately considered as binary. Concretely, we developed an ID estimator for binary data, called BID (from Binary Intrinsic Dimension) which, as we show below, can cope numerically with extremely large dimensionalities. Binary variables are the basis of all digital computation but are also of fundamental importance in both theoretical and applied science. The so-called spin systems, in which a spin is a binary degree of freedom, provide standard models to understand phases of matter, such as spin glassesBinder and Young [1986], and spin liquidsSavary and Balents [2016], and the corresponding phase transitions, like spontaneous orderingKardar [2007]. Binary variables are also used to describe biological neuronsHopfield [1982], Engel and Broeck [2001] or social dynamicsCastellano et al. [2009]. In artificial neural networks, the activations are typically real numbers, but many efforts focus on finding more efficient representations: the so-called quantization methodsHubara et al. [2018], Guo [2018] aim to represent activations or weights using small integer numbers or binary variables, in order to drastically save memory, energy, and computing time, while preserving performance. Activations and weights can be transformed in binary variables by the sign functionCourbariaux et al. [2016], Wang et al. [2023]. Binarization has been shown to preserve the scalar product between the activations and weightsAnderson and Berg [2017]. In Supp. Inf. we show that binarization also approximately preserves the local neighbourhoods computed using the full-precision activations inside a Large Language Model (LLM). This suggests that binarizing in high dimensions is an operation that retains relevant information about the dataGlielmo et al. [2022a].
@@
$P_{0}(r)=\frac{1}{2^{N}}\binom{N}{r}.$ (1)
Geometrically, the coordinates of uncorrelated random bits are points uniformly distributed in the $N$-dimensional configuration space $\mathcal{C}$, and the intrinsic dimension $d$ coincides with the number of variables, $N$. If one would measure the empirical distribution of the hamming distance, by following a procedure akin to the one used in ref.Macocco et al. [2023], one could infer the value of $N$ by the shape of such a distribution. However, such a procedure would fail if the bits are correlated since the empirical distribution would not match the theoretical distribution in eq. 1 for any value of $N$. In the case of correlated spins, we assume that the dimension is a smooth function of the distance $r$, and therefore make the following ansatz:
@@
where $C$ is the normalization constant. Eq. (2) reduces to (1) for $d(r)=N$ and $C=1$. We empirically found that model (2) fits accurately the observed distributions, at least locally, if one retains the first two terms of the Taylor expansion:
@@
where $r^{*}$ is a meta-parameter that allows constraining the fit to small distances (see Supp. Inf.). As a representative measure of the BID of a system, we consider the limit for small distances of $d(r)$, which qualitatively corresponds to the local number of degrees of freedom. According to Eq. (3), we simply have $BID\equiv d_{0}$.
@@
We first show that our algorithm is able to compute intrinsic dimensions that scale linearly with size when the input data consists of statistically independent subsystems. We generated 2500 bit streams by concatenating $B$ blocks of $N=10^{4}$ bits, where each block is generated by independently harvesting spins from a Boltzmann distribution at $T=2$, with an energy function $E(\boldsymbol{\sigma})=\sum_{i=1}^{N=1}\sigma_{i}\sigma_{i+1}$ with periodic boundary conditions ($\sigma_{N+1}\equiv\sigma_{1}$). Table 1 shows that the BID per bit is approximately constant from $10^{4}$ bits to $10^{6}$ bits, consistent with the system being constructed by concatenating independent blocks. Remarkably, the approach allows estimating intrinsic dimensions of order $10^{5}$ using only $2500$ data points. This is empirical evidence that the estimator proposed in this work is not affected by the curse of dimensionality, and can be used in extremely high dimensional spaces.
@@
Figure 2: The BID of image-crop representations in Resnet18. Upper row of images: example of crops for a sample of ImageNet. On top of each image, we report the size of the crop. All images are resized to $224\times 224$ resolution after cropping, thus the rightmost image is the full image. Lower row of images: a set of images constructed from random patches of size $28\times 28$ from ImageNet samples. A, B, C, D panels: BID of image representations as a function of the number of pixels in the cropped image, $N_{crop}$, where we normalized the latter by dividing by the total number of pixels in the full image, $N_{tot}=224^{2}$. For each panel, $l$ is the layer index and $L_{R}$ is the total number of layers. For further details see Supp. Inf.
@@
As a last example, we show that the BID can be used to detect correlations in the feature maps learned by large transformer-based architectures trained on a natural language processing (NLP) task. The ID has been used to analyze the properties of deep autoregressive transformers for protein sequence analysis and image analysisValeriani et al. [2023] and more recently for language Cheng et al. [2023, 2024]. However, to make the ID estimation possible in Ref. Valeriani et al. [2023] the representations were averaged in the sequence dimension, mixing information associated with the different tokens. In Refs. Cheng et al. [2023, 2024], the analyses are performed only on the representation of the last token. We here use our approach to estimate the binary intrinsic dimension of the representation of whole sentences, probing in this manner the collective properties of all the neurons of a layer. We consider paragraphs from Wikitext consisting of more than $300$ tokens, and we study how the BID scales with text length, binarizing activations through the application of the sign function. The initial representation (layer 0, Fig. 3, left panel) has short-range correlations, signaled by the BID per bit slowly drifting up as a function of the number of tokens (Fig. 3). Token embeddings are correlated because they must be compliant with the grammar structure of natural language, in which words are often used in groups. For long sentences, the BID scales approximately linearly with the sentence length: grammar-related correlations become less relevant in parts of a sentence separated by hundreds of words. Instead, the BID per bit of the last-layer representation learned by the model displays a clear power-law decay as a function of the number of tokens, similar to the one observed in scale-free critical spin systems (see Fig. 1A). Since the semantic content in text is a collective property of the whole sentence, the network to capture this property develops a long-range correlated representation: qualitatively the process of understanding the meaning of a long sentence can be described as the process of finding meaningful correlations in the sequence of tokens. To validate that we are indeed capturing the semantic correlations in text through the BID scaling we repeat the experiment by concatenating 150 tokens from two randomly chosen sentences from the same corpus. We observe that the power law decrease stops precisely at the boundary between phrases, and it is followed by an approximate plateau, consistent with the two halves of the paragraph containing independent information. In the last last part of the plot, the BID per bit starts decreasing again, as a result of the semantic correlations of the second half of the sentence. In Supp. Inf. we show that these results hold also for changing both corpus and network.
@@
In conclusion, we introduced an intrinsic dimension estimator specifically designed for binary data. To the best of our knowledge, the only other estimator for such variables was recently introduced in the context of formal concept analysis Hanika and Hille [2024]. However, it has only been benchmarked on simple binary data tables, whereas our approach is tailored for large bit streams. We tested examples with up to $10^{6}$ bits without observing systematic errors in the scaling of the estimation with system size.
@@
It is noteworthy that the last two datasets consist of real-valued features, while the BID estimator is designed for binary data. Each feature was converted into a binary variable by taking its sign. In Supp. Inf. we show that the trends observed in Fig. 3 hold when using two bits per feature instead of one. Specifically, the relative BID estimated for this higher-precision representation scales with the same power law. This suggests that when the intrinsic dimension becomes extremely large (around 1000 or more), its trends are insensitive to the precision used to represent individual features, and binning retains the essential information of the representation, as suggested in Refs. Hubara et al. [2018], Courbariaux et al. [2016].
The BID of data representations is clearly not restricted to either images or text, which are taken here as use cases where semantic correlations are easily interpretable. We expect our algorithm to be useful in the analysis of generic high dimensional spaces where global correlations are unknown, for example data coming from stock markets or brain activity. In perspective it would be interesting to explore its relationship with the Shannon entropy estimated by compression algorithmsAvinery et al. [2019].

### v2

@@ Abstract
Correlations among the features defining a data point force a data set to live in a manifold $\mathcal{M}$ whose geometry encodes the correlation structure of data. The most fundamental property of $\mathcal{M}$ is its intrinsic dimensionBennett [1969] (ID) which is qualitatively defined by the number of independent variables needed to parameterize it without significant information loss.
If one was able to estimate the intrinsic dimension in a generic big data scenario, one would be able to quantify the strength of correlations in real-world datasets, like a corpus of sentences or a large set of measurements from a camera. However, real-world data are represented in feature spaces that are outrageously large, and most of the available methods meet serious problems if the intrinsic dimension of the manifold is of the order of hundreds or moreCamastra and Staiano [2016]. We developed an intrinsic dimension estimator that generalizes the concept of the intrinsic dimension to the case of binary variables. We show that this allows circumventing the curse of dimensionality, estimating reliably the intrinsic dimension even when it is of order 100,000. This result paves the way to a plethora of applications, allowing to address questions which have thus far been, to the best of our knowledge, intractable. For example, we can measure the range of semantic correlations in a corpus of sentences with hundreds of words, and we demonstrate that these correlations are destroyed when shorter, unrelated sentences are joined together. We show that our approach is effective in a variety of other fields and topics, from the study of phase transitions and of frustration-induced disorder in statistical physics, to image classification.
Detecting and quantifying correlations in generic real-world data is a complex task since features live in high dimensional spaces and have interactions of arbitrary range and nature. For example, in an image, the meaning of a scene can emerge from the simultaneous presence of two objects that are physically far away. Moreover, semantic correlations are often associated with collective effects, such as the simultaneous presence of three or more specific objects. A prominent example where non-trivial correlations emerge is the feature space of deep neural networks. Those spaces typically lack a well-defined order, and neurons that are close in the architecture are not necessarily coding related information. However, the activations are known to be contained in a manifold of relatively low dimensionGoldt et al. [2020], and the existence of this manifold is believed to be an important ingredient in the learning process.
@@
In order to estimate the ID in very large feature spaces, we here renounce to develop a general-purpose ID estimator, and we focus on features that are binary or can be approximately considered as binary. Concretely, we developed an ID estimator for binary data, called BID (from Binary Intrinsic Dimension) which, as we show below, can cope numerically with extremely large dimensionalities. Binary variables are the basis of all digital computation but are also of fundamental importance in both theoretical and applied science. The so-called spin systems, in which a spin is a binary degree of freedom, provide standard models to understand phases of matter, such as spin glassesBinder and Young [1986], and spin liquidsSavary and Balents [2016], and the corresponding phase transitions, like spontaneous orderingKardar [2007]. Binary variables are also used to model social dynamicsCastellano et al. [2009], biological neuronsHopfield [1982] and learning processesEngel and Broeck [2001], Barbier et al. [2025], Marino and Ricci-Tersenghi [2024]. In artificial neural networks, the activations are typically real numbers, but many efforts focus on finding more efficient representations: the so-called quantization methodsHubara et al. [2018], Guo [2018] aim to represent activations or weights using small integer numbers or binary variables, in order to drastically save memory, energy, and computing time, while preserving performance. Activations and weights can be transformed in binary variables by the sign functionHubara et al. [2016], Wang et al. [2023]. Binarization has been shown to preserve the scalar product between the activations and weightsAnderson and Berg [2018]. In Supp. Inf. we show that binarization also approximately preserves the local neighbourhoods computed using the full-precision activations inside a Large Language Model (LLM). This suggests that binarizing in high dimensions is an operation that retains relevant information about the dataGlielmo et al. [2022a].
@@ I Binomial estimator
$P_{0}(r)=\binom{N}{r}p^{r}(1-p)^{N-r},$ (1)
with $p=1/2$, which can take the compact form $P_{0}(r)=\frac{1}{2^{N}}\binom{N}{r}$. Geometrically, the coordinates of uncorrelated random bits are points uniformly distributed in the $N$-dimensional configuration space $\mathcal{C}$, the corners of a high dimensional hypercube, and the intrinsic dimension $d$ coincides with the number of variables, $N$. In the case of correlated spins, we assume that the dimension is a smooth function of the distance $r$, and therefore make the following ansatz:
@@
where $C$ is a normalization constant. Eq. (2) reduces to (1) for $d(r)=N$ and $C=1$. We empirically found that model (2) fits accurately the observed distributions, at least locally, if one retains the first two terms of the Taylor expansion:
@@
where $r^{*}$ is a meta-parameter that allows constraining the fit to small distances (see Supp. Inf.). As a representative measure of the BID of a system, we consider the limit for small distances of $d(r)$, which qualitatively corresponds to the local number of degrees of freedom. This definition is consistent with the idea that, if the data are defined by real variables, the ID should be equal to the dimension of the hyperplane that is tangent to the data manifold, which can be estimated considering only small distances. According to Eq. (3), we simply have $BID\equiv d_{0}$. The computational complexity of computing our observable $P_{emp}(r)$ scales linearly with the number of features and quadratically with the number of samples, since it involves computing all the Hamming distances between each pair of points.
Eq. (4) allows us to quantify rigorously how well our model approximates the observed data, $P_{emp}(r)$. In order to assign an uncertainty to the value obtained by the optimization, one can perform Bayesian inference with a Markov Chain Monte Carlo (MCMC) on the parameter space of model (2), given the observations. In Supp. Inf. we show that both the minimization of Eq. (4) and MCMC Bayesian inference with a uniform prior provide similar results. MCMC allows estimating the posterior distribution of two parameters of the model, and, as a consequence their mean and variances.
II Results
@@
We first show that our algorithm is able to compute intrinsic dimensions that scale linearly with size when the input data consists of statistically independent subsystems. We generated 2500 bit streams by concatenating $B$ blocks of $N=10^{4}$ bits, where each block is generated by independently harvesting spins from a Boltzmann distribution at $T=2$, with an energy function $E(\boldsymbol{\sigma})=\sum_{i=1}^{N=1}\sigma_{i}\sigma_{i+1}$ with periodic boundary conditions ($\sigma_{N+1}\equiv\sigma_{1}$). Table 1 shows that the BID per bit is approximately constant from $10^{4}$ bits to $10^{6}$ bits, consistent with the system being constructed by concatenating independent blocks. Remarkably, the approach allows estimating intrinsic dimensions of order $10^{5}$ using only $2500$ data points. This is empirical evidence that the estimator proposed in this work is not affected by the curse of dimensionality, and can be used in extremely high dimensional spaces. For a detailed analysis of the dependence on the number of samples see Sec. A.2 of Supp. Inf.
@@
Figure 2: The BID of image-crop representations in Resnet18. Upper row of images: example of crops for a sample of ImageNet. On top of each image, we report the size of the crop. All images are resized to $224\times 224$ resolution after cropping, thus the rightmost image is the full image. Note moreover that the crop is performed only along the spatial dimensions, keeping the three channels of the images. Lower row of images: a set of images constructed from random patches of size $28\times 28$ from ImageNet samples. A, B, C, D panels: BID of image representations as a function of the number of pixels in the cropped image, $N_{crop}$, where we normalized the latter by dividing by the total number of pixels in the full image, $N_{tot}=224^{2}$. For each panel, $l$ is the layer index and $L_{R}$ is the total number of layers. For further details see Supp. Inf.
@@
As a last example, we show that the BID can be used to detect correlations in the feature maps learned by large transformer-based architectures trained on a natural language processing (NLP) task. The ID has been used to analyze the properties of deep autoregressive transformers for protein sequence analysis and image analysisValeriani et al. [2023] and more recently for language Cheng et al. [2023, 2025]. However, to make the ID estimation possible in Ref. Valeriani et al. [2023] the representations were averaged in the sequence dimension, mixing information associated with the different tokens. In Refs. Cheng et al. [2023, 2025], the analyses are performed only on the representation of the last token. We here use our approach to estimate the binary intrinsic dimension of the representation of whole sentences, probing in this manner the collective properties of all the neurons of a layer. We consider paragraphs from Wikitext consisting of more than $300$ tokens, and we study how the BID scales with text length, binarizing activations through the application of the sign function. The initial representation (layer 0, Fig. 3, left panel) has short-range correlations, signaled by the BID per bit slowly drifting up as a function of the number of tokens (Fig. 3). Token embeddings are correlated because they must be compliant with the grammar structure of natural language, in which words are often used in groups. For long sentences, the BID scales approximately linearly with the sentence length: grammar-related correlations become less relevant in parts of a sentence separated by hundreds of words. Instead, the BID per bit of the last-layer representation learned by the model displays a clear power-law decay as a function of the number of tokens, similar to the one observed in scale-free critical spin systems (see Fig. 1A). Since the semantic content in text is a collective property of the whole sentence, the network to capture this property develops a long-range correlated representation: qualitatively the process of understanding the meaning of a long sentence can be described as the process of finding meaningful correlations in the sequence of tokens. To validate that we are indeed capturing the semantic correlations in text through the BID scaling we repeat the experiment by concatenating 150 tokens from two randomly chosen sentences from the same corpus. We observe that the power law decrease stops precisely at the boundary between phrases, and it is followed by an approximate plateau, consistent with the two halves of the paragraph containing independent information. In the last part of the plot, the BID per bit starts decreasing again, as a result of the semantic correlations of the second half of the sentence. In Supp. Inf. we show that these results hold also for changing both corpus and network.
@@
III Comparison with other methods
Figure 4: Comparison between BID and other real-space estimators. Left panel: thermal dependence of the intrinsic dimension for a small $30\times 30$ ferromagnetic Ising model. Right panel: scaling of the intrinsic dimension normalized per bit at temperature $T=1.9$. Number of samples: 5000.
We compared our results with those obtained with other intrinsic dimension estimators, in particular: (i) the Correlation DimensionGrassberger and Procaccia [1983] (CD), perhaps the simplest and most widely know approach for estimating the ID; (ii) the MLE estimator in Ref. Levina and Bickel [2004] and GRIDEDenti et al. [2022], which estimate the ID by maximizing the likelihood of the empirical probability of the ratio of the distances of the nearest neighbors. (iii) I3DMacocco et al. [2023], which is designed to estimate the intrinsic dimension in data spaces in which the features take discrete values, and (iv) FCIErba et al. [2019], an ID estimator that was shown to infer the correct number of dimensions for up to 30 non-interacting spins and also for toy distributions in dimensions of the order of hundreds.
First, for a small 30x30 ferromagnetic Ising model we studied the thermal dependence of the ID (Fig. 4, left). At low temperature the system is in the ferromagnetic phase, which implies that a large fraction of the spins points in the same direction. We observe for all estimators except FCI that in these conditions the estimated ID is small and of roughly comparable magnitude. The estimates are consistent because in this special regime the local estimators (MLE, GRIDE, I3D and CD) are not severely affected by the dimensionality. In the same regime, the FCI estimator predicts a qualitatively different behavior, with the ID being very high and a decreasing function of T. This is at odds with physical intuition. Indeed, in a Ising model at low temperature only few defects are present, and the state of the system can be fully described by providing only the position of those defects. Therefore, the ID should smoothly tend to zero for $T\rightarrow 0$, similarly to the entropy. We notice that in the ferromagnetic phase most of the spins have the same value. In these conditions the centering and normalization procedure at the basis of FCI places the data on a hypersphere, but the distribution of the data on it is highly non-uniform. The qualitatively different results obtained by FCI might then be induced by the violation of one of its assumptions, that the data are isotropic when observed from the geometric center of the data set. In the paramagnetic phase, the BID estimate is instead comparable to the FCI estimate. Indeed, in this phase the spins are not polarized, namely the magnetization defined as $M=\sum_{i=1}^{N}\sigma_{i}$ is null. In these conditions, after centering and normalization, the data will be approximately isotropic, making the estimate by FCI more reliable. In contrast, all the other local estimators subestimate the value of the ID because they need an exponential number of samples in order to converge.
The right panel of Fig. 4 shows the size scaling of the different algorithms at fixed low temperature, where we know the system is extensive, and thus the ID must scale linearly with the number of bits, $N$. For small system sizes we observe that the local estimators (MLE, GRIDE, I3D and CD) give results comparable with the BID, but their estimate of the ID does not scale linearly with $N$, at odds with physical evidence. This is the behavior which prompted us to develop a new estimator, specifically designed for bits or spins. The BID scales linearly with the number of bits, and so does FCI. Nonetheless for FCI the slope of the line is one order of magnitude bigger and, as discussed above, the corresponding estimation is also at odds with physical intuition.
We finally performed ID estimates with GRIDE and FCI on the images and text databases. For our experiment on images, we observe that both GRIDE and FCI behave qualitatively similarly to the BID, namely the plateau observed for real data and its absence for images constructed grouping random patches are reproduced (see Sec. A.8.1 of Supp. Inf.). However, we remark that the ID estimation performed with the GRIDE is more than one order of magnitude smaller than the estimations done by BID, whereas the estimations of FCI are systematically below that of BID, by roughly a factor of 2. Additionally, in Sec. A.8.3 of Supp. Inf. we study the BID of simple images directly in pixel space, for MNIST and CIFAR10. There we show that the BID of MNIST is fully compatible with other real-space estimators, and we discuss the challenges of finding correct binarizations for RGB images. For the text experiment, we show in Sec. A.8.2 of Supp. Inf. that GRIDE fails to display a plateau at the semantic boundary since, as we showed in Fig. 4, it doesn’t scale correctly with system size. Instead, the FCI estimator produces qualitatively comparable results to ours, reproducing the linear scaling for the ID at the layer zero and detecting the semantic boundary in the case of concatenated paragraphs. Nonetheless, we observed significant differences in the size scaling exponents of the BID normalized per bit at the last layer of the transformer, namely the slope $\eta$ of the lines in the log-log plot of $BID(T)/N\propto T^{-\eta}$. Whereas for the BID we get an exponent of $\eta_{BID}=0.42$, for FCI and GRIDE we get much higher values, $\eta_{FCI}=0.89$, and $\eta_{GRIDE}=0.70$. The study of the relationship between the BID and the scaling exponents in systems with power-law correlations, and in particular in critical systems, together with the quantitative discrepancies observed, are interesting topics which we are currently investigating.
Wrapping up, if the ID is small, our estimator is qualitatively consistent with all the other estimators we considered, except FCI if the bits have a preferential value (like in a ferromagnetic phase). When the ID is large, our estimator provides values closer to the FCI estimator, while the other estimators provide lower values. Only our estimator and the FCI estimator provide ID estimates which scale linearly with system size, although in Fig. 4 we observed that the FCI values of the ID per spin are almost one order of magnitude larger in the ordered phase than in the disordered phase, a result that is not plausible.
The significant quantitative differences with respect to MLE, CD, I3D, GRIDE and FCI observed in some regimes can be understood by considering that the estimate of the ID depends on the assumption on the geometrical structure of the manifold containing the data. For binary data this manifold is a N-dimensional hypercube, and the data can only occupy its vertices. The I3D estimator assumes that the manifold is an ID-dimensional cubic lattice. The resulting estimator correctly captures the dimension of points with discrete-valued features when the intrinsic dimension is moderate. However, the difference in the geometric structure of the manifold (a cube of size one for BID, a lattice of infinite size for I3D) determines qualitative differences when the ID is large. The CD, MLE, GRIDE and FCI are all implicitly or explicitly based on the assumption that the features are real numbers. In FCI it is also explicitly assumed that the data are isotropic, namely that that are uniformly present in all the directions with respect to the geometric center of the dataset. These assumptions do not hold, even approximately, in binary data, bringing to systematic errors.
IV Conclusions
@@
We introduced an intrinsic dimension estimator specifically designed for binary data. To the best of our knowledge, the only other estimator for such variables was recently introduced in the context of formal concept analysis Hanika and Hille [2024]. However, it has only been benchmarked on simple binary data tables, whereas our approach is tailored for large bit streams. We tested examples with up to $10^{6}$ bits without observing systematic errors in the scaling of the estimation with system size.
@@
It is noteworthy that the last two datasets consist of real-valued features, while the BID estimator is designed for binary data. Each feature was converted into a binary variable by taking its sign. In Supp. Inf. we show that the trends observed in Fig. 3 hold when using two bits per feature instead of one. Specifically, the relative BID estimated for this higher-precision representation scales with the same power law. This suggests that when the intrinsic dimension becomes extremely large (around 1000 or more), its trends are insensitive to the precision used to represent individual features, and binning retains the essential information of the representation, as suggested in Refs. Hubara et al. [2018, 2016].
The BID of data representations is clearly not restricted to either images or text, which are taken here as use cases where semantic correlations are easily interpretable. We expect our algorithm to be useful in the analysis of generic high dimensional spaces where global correlations are unknown, for example data coming from stock markets or brain activity. There are a number of interesting perspectives, among which we can mention understanding the influence of different binarization protocols of generic continuos data on the resulting BID; the use of analytical solutions in simple models for the distribution of Hamming distances, $P(r)$, to gain additional physical insights on the respective BID; and exploring the relationship between the BID and the Shannon entropy estimated by compression algorithmsAvinery et al. [2019].
Code availability
Data availability
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.